# HiCoLoRA: Addressing Context-Prompt Misalignment via Hierarchical Collaborative LoRA for Zero-Shot DST

### v2

@@ 5 Conclusion
zs-DST is crucial for scalable TODs but remains challenged by insufficient cross-layer coordination, semantic conflation across domains, and corruption of pre-trained knowledge. HiCoLoRA overcomes these issues via a hierarchical LoRA design for dynamic context-prompt alignment, spectral clustering for domain-slot disentanglement, and SemSVD-Init for knowledge-preserving fine-tuning. Evaluations on MultiWOZ and SGD show that HiCoLoRA significantly outperforms previous SOTA approaches, improving average JGA by 5.4% and 9.4%, respectively. Limitations remain in highly idiosyncratic slot domains; future work will focus on slot-aware refinement to further strengthen HiCoLoRA’s applicability in zs-DST.
Appendix A Additional Experimental Results
A.1 Performance on SGD Dataset
SlotDM-DST 2022 43.9/86.3 36.6/61.4 16.5/62.0 46.7/86.9
T5DST 2021 46.8/N 48.8/N 55.5/N 59.2/N 23.3/N 53.0/N
Prompter 2023 48.4/N 51.5/N 65.3/N 59.2/N 21.9/N 50.8/N
DCC 2023 28.8/N 19.4/N 42.3/N
Table 2 presents the zero-shot performance of HiCoLoRA on the SGD Dataset. Compared to both baseline methods and previous state-of-the-artapproaches, HiCoLoRA achieves significant improvements across multiple domains.
A.2 Comparison with Contemporary LoRA Methods
Method Year Attr. Hotel Train Taxi Rest. AVG.
HydraLoRA 2024 35.1 18.9 26.3 41.5 65.2 37.4
LoRA-GA 2024 33.8 19.2 24.7 42.8 64.1 36.9
RoSA 2024 36.5 19.6 27.9 43.2 66.8 38.8
Spectral Adapter 2025 37.2 20.1 28.5 43.6 67.3 39.3
HiCoLoRA (Ours) 2025 38.9 20.4 31.0 44.9 68.6 40.8
To situate HiCoLoRA within the evolving landscape of PEFT methods, we compare it against four contemporary LoRA variants: HydraLoRA Tian et al. (2024), LoRA-GA Wang et al. (2024b), RoSA Nikdan et al. (2024), and Spectral Adapter Zhang & Pilanci (2025). As shown in Table 3, HiCoLoRA achieves the highest average JGA, outperforming all baselines across nearly all domains. This superiority is not merely incremental; it stems from fundamental architectural and semantic distinctions that address core challenges in zs-DST.
Knowledge Preservation and Transfer: LoRA-GA improves gradient alignment during initialization to accelerate convergence, but it does not explicitly modulate singular values to align with task-specific semantics. HiCoLoRA’s SemSVD-Init not only preserves pre-trained knowledge but also amplifies singular components relevant to domain-slot structures, effectively mitigating catastrophic forgetting and enhancing zero-shot generalization, particularly for rare slots such as hotel-stars.
HiCoLoRA address the unique challenges of zs-DST: cross-layer misalignment, semantic conflation, and knowledge distortion. While other LoRA variants offer general-purpose efficiency, HiCoLoRA provides a domain-aware and layer-conscious design that is essential for robust zero-shot transfer in TODs.
A.3 Scalability Analysis: Generalization Across Model Scales
Method Year Base Model Attr. Hotel Train Taxi Rest. AVG.
ChatGPT-zsTOD 2023 ChatGPT (GPT-3.5) 52.7 42.0 60.8 70.9 55.8 56.4
ChatGPT-zsTOD 2023 ChatGPT (GPT-3.5) 67.2 37.6 67.3 74.4 60.1 61.3
D0T 2024 LLAMA2-13B 63.1 43.8 60.8 48.8 64.7 56.2
MoPE 2024 ChatGLM-6B 60.4 34.1 64.0 71.3 55.9 57.1
FnCTOD 2024 ChatGPT (GPT-4) 58.8 45.2 69.5 76.4 63.2 62.6
FnCTOD 2024 LLAMA2-13B 62.2 46.8 60.9 67.5 60.3 59.5
Multi-User 2025 GPT-4o 56.8 46.0 61.9 69.3 55.1 57.8
HiCoLoRA 2025 LLAMA2-13B 62.0 42.0 61.0 65.0 69.0 60.0
HiCoLoRA 2025 Qwen2.5-14B-Instruct 64.0 44.0 63.0 68.0 71.0 62.0
To rigorously assess the scalability and architectural generality of HiCoLoRA, we extend our evaluation to LLM, comparing against contemporary LLM-based zs-DST methods, including ChatGPT-zsTOD Heck et al. (2023), D0T Finch & Choi (2024), MoPE Tang et al. (2024), FnCTOD Li et al. (2024) and Multi-User Song et al. (2025). As shown in Table 4, HiCoLoRA achieves competitive performance when deployed in LLAMA2-13B and Qwen2.5-14B-Instruct, with an average JGA of 62.0% in the latter, only marginally below FnCTOD with GPT-4 (62.6%) and significantly outperforms other baselines based on LLM.
Architectural Generalization Beyond Scale. The consistent performance of HiCoLoRA in both both small (T5-small, 60M) and large (13B–14B) models underscores a key insight: its hierarchical adaptation mechanism is scale-agnostic. The efficacy of HiCoLoRA stems from its structured semantic alignment decomposition, which addresses cross-layer coordination (RQ1), domain-slot disentanglement (RQ2), and knowledge preservation (RQ3) through explicit inductive biases. This allows it to be generalized effectively even when applied to larger models without architecture-specific modifications.
Limitations and Future Directions. The remaining gap between the HiCoLoRA and GPT-4-based methods suggests that scale still matters to capture extremely nuanced or idiosyncratic slot semantics. However, HiCoLoRA’s strong performance in structured domains such as Restaurant indicates that its hierarchical and spectral mechanisms effectively compensate for scale limitations through better semantic organization. Future work may explore hybrid approaches that integrate HiCoLoRA’s alignment mechanisms with larger foundation models for even stronger zero-shot generalization.
A.4 Ablation Study Table
Method Attr. Hotel Train Taxi Rest. AVG.
HiCoLoRA (Full) 38.9 20.4 31.0 44.9 68.6 40.8
w/ Swap Hier Strategies 37.2 19.7 22.9 40.2 67.5 37.4
w/o Adaptive Fusion 28.9 19.3 20.3 43.0 68.0 35.9
w/o Spec Joint Cluster 36.2 19.8 27.5 42.1 63.6 37.8
w/ Kiming Init 34.3 20.4 27.8 40.4 67.5 38.1
w/ PiSSA Init 36.5 20.3 29.0 42.5 67.8 38.9
w/ MiLoRA Init 34.1 19.9 26.2 38.5 62.9 36.3
Table 5: Ablation study on hierarchical architecture, adaptive fusion, spectral clustering, initialization of HiCoLoRA on MultiWOZ. Attr. and Rest. are abbreviations for Attraction and Restaurant, respectively.
Table 5 validates the contributions and necessity of each core component of HiCoLoRA to its overall performance. This validation is conducted by systematically removing or replacing the core components, including the hierarchical strategy, adaptive fusion, spectral clustering, and initialization method.
Appendix B Experiments Setting Details
B.1 Dataset Statistic
@@
Based on the experimental design for zero-shot dialog state tracking, domain selection was strategically constrained to ensure robust evaluation. For MultiWOZ (Table 7), the police (46 dialogs) and hospital (38 dialogs) domains were excluded due to insufficient dialog volume and slot diversity, which would compromise statistical reliability in zero-shot generalization tests. Similarly, in SGD (Table 7), services with limited samples or atypical slot structures, such as RideSharing (Test: 112), Calendar (Test: 98), etc., are omitted to avoid skewing results. This curation focuses evaluation on domains with adequate data density and representative slot semantics, ensuring that performance metrics reflect true zero-shot transferability rather than data-sparsity artifacts. Consequently, while coverage is reduced, the core challenge of cross-domain adaptation is preserved, with results generalizable to mainstream service-oriented interactions.
B.2 Baseline Models
@@
B.2.1 Main Baseline
  • TRADE Wu et al. (2019) enhances dialog state generation by incorporating a copy mechanism and enabling cross-task knowledge transfer, allowing the model to handle unseen dialog states during training.
@@ B.2.1 Main Baseline
  • SUMBT Lee et al. (2019), built on BERT-base, employs contextual semantic attention to learn domain–slot-type and slot-value relations, predicting slot values in a non-parametric manner.
@@
  • SlotDM-DST Wang et al. (2022), leveraging T5-small, models slot–slot, slot–value, and slot–context dependencies via slot prompts, value demonstrations, and constraint objects. Shared prompts capture transferable knowledge across domains.
@@
  • DualLoRA Luo et al. (2024) builds upon PPTOD-small with a T5-small backbone, employing two low-rank adaptation matrices, one refining dialog context and the other slot prompts. Once trained, these matrices are fused into the frozen pre-trained weights, yielding zero-shot cross-domain dialog state tracking without any extra inference latency.
B.2.2 LoRA Baseline
  • HydraLoRA Tian et al. (2024) is a parameter-efficient fine-tuning (PEFT) framework designed to address the performance gap between standard LoRA and full fine-tuning, especially on complex datasets. It introduces an asymmetric LoRA structure that requires no domain expertise. Experiments demonstrate that HydraLoRA surpasses existing PEFT methods in performance.
  • LoRA-GA Wang et al. (2024b) enhances LoRA by proposing a novel gradient-aware initialization strategy that aligns the gradients of the low-rank matrices with those of full fine-tuning at the first training step. This method significantly accelerates convergence (2–4× faster than vanilla LoRA) and improves performance across tasks such as GLUE, GSM8K, and code generation, even for large models like Llama 2-7B.
@@ B.2.2 LoRA Baseline
  • Spectral Adapter Zhang & Pilanci (2025) incorporates spectral information from pretrained weights via SVD to enhance PEFT methods. It performs additive tuning or orthogonal rotation on the top singular vectors, improving rank capacity and parameter efficiency. The adapter also benefits multi-adapter fusion and demonstrates stronger performance across various tasks.
B.2.3 LLM Baseline
@@
  • D0T Finch & Choi (2024) enhances zero-shot DST by generating synthetic data across over 1,000 domains, creating a diverse training dataset with silver-standard annotations. This approach addresses data scarcity and enables adaptation to new domains without costly collection efforts.
@@
  • FnCTOD Li et al. (2024) improves zero-shot DST through function calling with LLMs, allowing adaptation to diverse domains without extensive data or tuning. It achieves state-of-the-art performance with both open-source and proprietary LLMs, significantly boosting ChatGPT and GPT-4 results.
  • Multi-User Song et al. (2025) assesses LLMs in multi-user DST by extending datasets with second-user utterances generated via speech act theory. For a fair comparison, the experimental setup was configured using single-user data to evaluate the performance of LLMs in single-user dialog state tracking.
B.3 Evaluation Metric Formulas
B.3.1 JGA Formula
@@ B.3.2 AGA Formula
In this formula, $T$ denotes the total number of dialog turns in the evaluation dataset. For each turn $i$, $S_{i}^{pre}$ and $S_{i}^{gt}$ represent the predicted and ground truth sets of slot-value pairs, respectively. The formula calculates slot-level accuracy for each turn by:
@@
  • Normalizing by the total number of ground truth slot-value pairs $|S{i}^{gt}|$
@@
B.4 Experiments Implementation Details
Our experimental setup, designed for precise comparison with prior work, follows that of DualLoRA Luo et al. (2024). We use the T5-small architecture (6 encoder/decoder layers, 512 hidden dimension, 8 attention heads) as the backbone for HiCoLoRA, with a LoRA rank of 8 for low-rank adaptation, initialized from PPTOD-small checkpoints, consistent with observations in DualLoRA that PPTOD Su et al. (2022) is particularly suitable for prompt-tuning due to its pre-training objectives.

### v3

@@ 6 Conclusion
zs-DST is crucial for scalable TODs but remains challenged by insufficient cross-layer coordination, semantic conflation across domains, and corruption of pre-trained knowledge. HiCoLoRA overcomes these issues via a hierarchical LoRA design for dynamic context-prompt alignment, spectral clustering for domain-slot disentanglement, and SemSVD-Init for knowledge-preserving fine-tuning. Evaluations in MultiWOZ and SGD show that HiCoLoRA significantly outperforms previous SOTA approaches, improving average JGA by 5.4% and 9.4%, respectively. Limitations remain in highly idiosyncratic slot domains, and future work will focus on slot aware refinement to further strengthen HiCoLoRA’s applicability in zs-DST.
Limitations
While HiCoLoRA advances zs-DST performance, several limitations persist: the model struggles with ambiguous slot boundaries, leading to prediction errors when slot values overlap or are implicitly referenced; cross domain confusion arises in multi-domain dialogs where similar slot names cause semantic entanglement; rare or unseen slot values are poorly generalized, as the current initialization and adaptation mechanisms do not fully address domain specific sparsity; and highly idiosyncratic slots with domain-exclusive terms remain challenging, as spectral clustering may fail to capture low frequency semantic associations and higher layer fusion can dilute relevant signals. Future work may refine slot boundary detection, enhance domain disambiguation, and develop more targeted rare slot handling strategies.
Ethics Statement
Our research involves only publicly available, anonymized dialog datasets (MultiWOZ and SGD) and does not collect new human subject data. Their use complies with the consent agreements established during their original release. All data usage complies with the original licenses, and no personally identifiable information is processed or stored. The proposed method, HiCoLoRA, is designed to improve zero-shot generalization in task-oriented dialog systems and does not have known harmful applications. While the method itself poses no direct ethical hazards, we note that any dialogue system carries a risk of misunderstanding user inputs or propagating biases if deployed without proper safeguards in sensitive applications. We acknowledge that there are no conflicts of interest and that the research was conducted with full integrity, transparency, and respect for privacy, fairness, and inclusivity. No institutional review board (IRB) approval was required as the study involves no human participants beyond the use of existing, de-identified benchmark data.
We acknowledge the use of Writeful integrated with Overleaf for refining the textual expression of this manuscript, and DeepSeek V3.2 for error correction of the experimental code. The role of these LLMs was limited to technical assistance and did not involve research ideation or the creation of core content. All LLM outputs have been rigorously verified by the authors, who bear full responsibility for the final accuracy, integrity, and originality of the content including the avoidance of plagiarism or scientific misconduct.
Appendix A Experimental Details
A.1 Dataset Statistic
@@
Based on the experimental design for zero-shot dialog state tracking, domain selection was strategically constrained to ensure robust evaluation. For MultiWOZ 2.1 (Eric et al., 2020) (Table 2), the Police (46 dialogs) and Hospital (38 dialogs) domains were excluded due to insufficient dialog volume and slot diversity, which would compromise statistical reliability in zero-shot generalization tests. Similarly, in SGD (Rastogi et al., 2020) (Table 3), services with limited samples or atypical slot structures, such as RideSharing (Test: 112), Calendar (Test: 98), etc., are omitted to avoid skew results. This curation focuses on evaluation on domains with adequate data density and representative slot semantics, ensuring that performance metrics reflect true zero-shot transferability rather than data-sparsity artifacts. Consequently, while coverage is reduced, the core challenge of cross-domain adaptation is preserved, with results generalizable to mainstream service-oriented interactions.
A.2 Baseline Models
@@
A.2.1 Main Baseline
  • TRADE Wu et al. (2019) enhances dialog state generation by incorporating a copy mechanism and enabling knowledge transfer between tasks, allowing the model to handle unseen dialog states during training.
@@ A.2.1 Main Baseline
  • SUMBT Lee et al. (2019), built on the BERT-base, employs contextual semantic attention to learn the domain-slot-type and slot value relations, predicting slot values in a non-parametric manner.
@@
  • SlotDM-DST Wang et al. (2022), leveraging T5-small, models slot-slot, slot-value, and slot-context dependencies via slot prompts, value demonstrations, and constraint objects. Shared prompts capture transferable knowledge across domains.
@@
  • DualLoRA Luo et al. (2024) builds on PPTOD-small with a T5-small backbone, employing two low-rank adaptation matrices, one refining dialog context and the other slot prompts. Once trained, these matrices are fused into the frozen pre-trained weights, yielding zero-shot cross-domain dialog state tracking without any extra inference latency.
A.2.2 LoRA Baseline
  • HydraLoRA Tian et al. (2024) is a parameter-efficient fine-tuning (PEFT) framework designed to address the performance gap between standard LoRA and full fine-tuning, especially on complex datasets. Introduce an asymmetric LoRA structure that does not require domain expertise. Experiments demonstrate that HydraLoRA surpasses existing PEFT methods in performance.
  • LoRA-GA Wang et al. (2024b) improves LoRA by proposing a novel gradient-aware initialization strategy that aligns the gradients of the low-rank matrices with those of full fine-tuning at the first training step. This method significantly accelerates convergence (2-4× faster than vanilla LoRA) and improves performance in tasks such as GLUE, GSM8K, and code generation, even for large models such as Llama 2-7B.
@@ A.2.2 LoRA Baseline
  • Spectral Adapter Zhang and Pilanci (2025) incorporates spectral information from pretrained weights via SVD to enhance PEFT methods. Performs additive tuning or orthogonal rotation on the top singular vectors, improving rank capacity and parameter efficiency. The adapter also benefits multi-adapter fusion and demonstrates stronger performance across various tasks.
A.2.3 LLM Baseline
@@
  • D0T Finch and Choi (2024) enhances zero-shot DST by generating synthetic data across over 1,000 domains, creating a diverse training dataset with silver-standard annotations. This approach addresses data scarcity and enables adaptation to new domains without costly collection efforts.
@@
  • FnCTOD Li et al. (2024) improves zero-shot DST by calling functions with LLMs, allowing adaptation to diverse domains without extensive data or tuning. It achieves state-of-the-art performance with both open-source and proprietary LLMs, significantly boosting ChatGPT and GPT-4 results.
  • Multi-User Song et al. (2025) evaluates LLMs in multi-user DST by extending datasets with second-user utterances generated via speech act theory. For a fair comparison, the experimental setup was configured using single-user data to evaluate the performance of LLMs in single-user dialog state tracking.
A.3 Evaluation Metric Formulas
A.3.1 JGA Formula
@@ A.3.2 AGA Formula
In this formula, $T$ denotes the total number of dialog turns in the evaluation dataset. For each turn $i$, $S_{i}^{pre}$ and $S_{i}^{gt}$ represent the predicted and ground truth sets of slot-value pairs, respectively. The formula calculates the slot-level accuracy for each turn by:
@@
  • Normalization by the total number of ground truth slot-value pairs $|S{i}^{gt}|$
@@
A.4 Experiments Implementation Details
Our experimental setup, designed for a precise comparison with previous work, follows that of DualLoRA Luo et al. (2024). We use the T5-small architecture (6 encoder/decoder layers, 512 hidden dimension, 8 attention heads) as the backbone for HiCoLoRA, with a LoRA rank of 8 for low-rank adaptation, initialized from PPTOD-small checkpoints, consistent with observations in DualLoRA that PPTOD Su et al. (2022) is particularly suitable for prompt-tuning due to its pre-training objectives.